OpenStudy (mitu12):

A toy company spends $1500 each day for factory expenses plus $8 per teddy bear. how many bears must the company sell in one day to equal it daily costs? write an equation and solve.

OpenStudy (anonymous):

can someone help me?

OpenStudy (mitu12):

with what

OpenStudy (anonymous):

with a question I posted earlier

OpenStudy (anonymous):

is right below this one

OpenStudy (dangerousjesse):

First of all; don't advertise your own questions on another persons thread. Second of all; don't post irrelevant replies in threads period.

OpenStudy (anonymous):

what? what do you mean?

OpenStudy (mitu12):

i need help to set up the equation

OpenStudy (dangerousjesse):

You're drawing attention away from someone else's thread for your own shiggles. Just post the link in chat.

OpenStudy (mitu12):

can someone help me to figure out the equation please

OpenStudy (dangerousjesse):

There isn't a price listed for the teddy bears, so I'm going to assume that they mean they profit 8 dollars per teddy bear.

OpenStudy (dangerousjesse):

So \(8x\geq 1500\)

OpenStudy (mitu12):

will the equation be 1500+8x=12x

OpenStudy (dangerousjesse):

Or if they mean 8 dollars per bear then x profit separately, \(x\times \textsf{total bears sold}\)\(\geq 1500+8\times\textsf{total bears bought}\)

OpenStudy (dangerousjesse):

Where did you get 12 from?

OpenStudy (mitu12):

ive seen some other questions with 12 in it

OpenStudy (dangerousjesse):

If there's no twelve in this question, then any other questions you've had are irrelevant..

OpenStudy (mitu12):

ohhh on the picture i saw a price of $12

OpenStudy (dangerousjesse):

I wish you would have said that earlier ;-;

OpenStudy (mitu12):

me too i just noticed that

OpenStudy (dangerousjesse):

Then 12x=8x+1500

OpenStudy (mitu12):

someone drew over it

OpenStudy (mitu12):

ok thank you so much

OpenStudy (mitu12):

375 bears total